# Adaptive Manipulation using Behavior Trees

### v2

@@ Abstract
Many manipulation tasks pose a challenge since they depend on non-visual environmental information that can only be determined after sustained physical interaction has already begun. This is particularly relevant for effort-sensitive, dynamics-dependent tasks such as tightening a valve. To perform these tasks safely and reliably, robots must be able to quickly adapt in response to unexpected changes during task execution. Humans can intuitively respond and adapt their manipulation strategy to suit such problems, but representing and implementing such behaviors for robots remains an open question. We present the adaptive behavior tree, which enables a robot to quickly adapt to both visual and non-visual observations during task execution, preempting task failure or switching to a different strategy based on data from previous attempts. We test our approach on a number of tasks commonly found in industrial settings. Our results demonstrate safety, robustness (100% success rate for all but one experiment) and efficiency in task completion (eg, an overall task speedup of 46% on average for valve tightening), and would reduce dependency on human supervision and intervention.
@@ I Introduction
Service robots can revolutionize how entire service-oriented domains work, with potential applications including autonomous maintenance in – often remote – industrial facilities [1, 2], such as oil rigs [3], or care robots in domestic environments [4, 5, 6]. To complete tasks effectively, robots must use observations from the environment to construct and execute a suitable plan; a number of frameworks have been developed for task and motion planning (TAMP) for manipulation [7, 8, 9, 10].
This work is motivated by effort-sensitive, dynamics-dependent manipulation tasks for articulated devices commonly found in industrial settings, for example twisting a valve until tightened to a specified torque or pushing an e-stop button until pressed. Robots generally require highly controlled environments with known task parameters for reliable planning and execution of manipulation tasks; however, the tasks we are interested in pose a significant challenge since they depend on information that can only be determined after sustained physical interaction has already begun. For example, a robot cannot determine the initial tightness of a valve or stiffness of a button from visual data alone, and can only gather this information by beginning to interact with it. Even then, in the absence of an accurate model for the dynamic profile for the device, the robot cannot know for how long it must twist the valve, or how hard it must push the button, to complete the task. Even if provided a model, stiction from corrosion or unclear servicing may lead to uncertainty and deviation from standard assumptions. To solve these kinds of tasks, we argue that the robot must be able to instead decide on an initial plan with which to start the task and adapt to observations it makes during task execution. Since the tasks we consider may require the robot to apply a significant amount of effort to manipulate the device, adaptability is not only critical for task completion but also for safe robot operation; failure to respond to a sudden unexpected spike in effort can cause catastrophic failure and even damage to the robot or environment.
@@
We also consider the utility of a robot having access to different manipulation strategies for the same task. A strategy is the specific approach that a robot may take to carry out a task. For manipulation tasks, different strategies may involve grasping the object differently, or even using a tool to assist manipulation. Different strategies can have different trade-offs; for example speed, complexity, effort and power expenditure. For example, quickly grasping and twisting a valve handle head-on may be optimal when beginning to twist an initially loose valve, but fail to provide the torque to finish tightening it, while grasping from the side to apply more wrench may be slower for most of the tightening procedure but can fully tighten the valve. It may even be that completely different strategies are required to successfully complete different task instances. We seek a system to coordinate this strategy selecting and switching behavior.
We tackle both problems by exploiting the reactivity and modularity of behavior trees. The key contribution of our work is the adaptive behavior tree, which enables a robot to handle uncertainty in the environment and quickly adapt to unexpected safety-critical changes during task execution, preempting task failure or switching to a more optimal strategy based on data gathered over previous attempts. It comprises two stages; a selector that uses data recorded over previous attempts to determine which strategy to deploy, and a monitor that checks a set of conditions concurrently to strategy execution, halting the current strategy if the task is complete or a safety condition is violated and re-attempting the task if necessary. The design is highly modular and readily allows for the addition and swapping out of strategy subtrees and condition nodes to suit a new task. We test our approach on various case studies representing complex, effort-intensive manipulation tasks for an industrial setting.
@@ II-A Adaptive Control for Robotics
It may be argued that our problem of manipulating articulated devices with unknown and changing dynamics intrinsics can be solved with an appropriate adaptive controller. Adaptive control has existed in the literature for many years, and generally focuses on online parameter estimation and learning for uncertain dynamic systems. There exist many good surveys on the literature for adaptive control [11, 12]. The approach has been extended to robotic manipulation tasks, for example estimating model parameters for an uncertain end-effector payload [13, 14, 15].
@@ II-B Behavior Trees for Robotics
There are many examples demonstrating how BTs can be used to coordinate manipulation tasks [20, 21, 22] and robot control [23, 24, 25]; Iovino et. al. provide a comprehensive survey of BTs for robotics [16]. BTs for robotics can be hand-coded [26, 27, 28], planned analytically [29, 30, 31, 32, 33, 34], or even learned from demonstration [35, 36].
@@ III Behavior Trees – An Overview
A Sequence propagates the tick to its children from left to right, returning Success if and only if all their children return Success. When it receives Running, it returns Running and ticks the same child on the next tick, and when it receives Failure, it returns Failure and resets the sequence progress. A ReactiveSequence behaves similarly, but when it receives Running, it returns Running and ticks all children up to the current child on the next tick. Thus, all prior children are still checked for Failure whenever it is ticked; this allows for reactive behavior.
A Fallback (known as a Selector in other works) propagates the tick to its children from left to right until one of them returns Success, at which point the Fallback stops and returns Success. When it receives Running, it returns Running and ticks the same child on the next tick. It returns Failure only if all children return Failure. A ReactiveFallback behaves similarly, but when it receives Running, it returns Running and ticks all children up to the current child on the next tick. Thus, all prior children are still checked for Success whenever it is ticked; this allows for reactive behavior much like in the case of a ReactiveSequence.
@@
A DecoratorNode modifies the return status of a single child node, for example forcing Failure. A Precondition is a special type of DecoratorNode that ticks the child only if a particular condition is met, or else returns either Success or Failure (marked by ‘S’ or ‘F’ respectively).
@@ IV Adaptive BT
Fig. 2 presents the proposed design for an adaptive BT, as well as an example of how it can be integrated within a wider robot system. The adaptive BT is composed of two core stages; the strategy selector first decides which strategy to use for the next attempt, based on current perception data as well as data recorded over past attempts (such as joint efforts over time). The condition monitor then wraps around the subtree for the selected strategy, checking a set of task- and strategy-specific conditions concurrently to strategy execution. The adaptive BT also admits subtrees to coordinate robot setup prior to strategy selection, and for safely exiting or recovering from the current attempt in preparation for the next.
Figure 2: Adaptive BT design, integrated within a wider robot system representing the test platform in this work. The coloured lines represent communication between the BT and the rest of the robot.
@@
The condition checking is such that if the subtree of task-specific conditions returns Success, the task is considered to have been completed successfully, and if it returns Failure, the task is still ongoing. This follows the BT design principle of backward chaining, commanding the robot to keep executing the strategy until the task completion conditions are met. Meanwhile, if the subtree of strategy-specific conditions returns Failure, a constraint has been violated during execution of the current strategy and the robot must halt the current attempt and try again, potentially selecting a different strategy. If it returns Success, no violations have occurred. This logic makes the condition monitor reactive and interpretable. If there are multiple task- or strategy-specific conditions to check, they should be placed in a ReactiveSequence to ensure proper operation of the adaptive BT. Each condition is ticked for every tick of the strategy subtree, allowing the robot to rapidly respond to new observations; thus the inherent reactivity of BTs is exploited. If a condition is not relevant to the current stage of the strategy subtree, we recommend skipping it by wrapping with an appropriate Precondition.
@@
The adaptive BT can easily be added as a subtree within a greater BT structure, and is thus readily applicable to existing BT-based robot task planning systems without major changes. The only required knowledge for adding the proposed design into a larger BT is that the adaptive BT must be repeatedly ticked while Running, and will return Success if the task was completed successfully and Failure otherwise. To avoid getting stuck, we recommend limiting the adaptive BT to a finite number of allowed attempts.
@@ V Adaptive BT for Manipulation
We now apply the proposed adaptive BT to the context of manipulation tasks wherein the robot must manipulate an articulated device; task success occurs when the device is considered sufficiently manipulated. For all manipulation tasks explored in this work, the manipulation procedure consists of identifying the pose of the device, then approaching, grasping, manipulating and finally retracting from the device. Fig. 3 presents the designs used for each subtree of the adaptive BT. We stress that these designs are just examples used to test the adaptive BT, with an emphasis on simplicity.
Figure 3: Subtrees used for all manipulation tasks considered in this work.
Setup. The robot moves to a pose suitable for observing the task instance to determine which strategy to select, and then estimates the pose of the device handle. If the device handle has rotational symmetry and thus has more than one valid pose, the robot later uses the most suitable pose that is within safe device handle pose limits for the selected strategy.
Strategy Subtree. The robot approaches, grasps and manipulates the device with its end-effector. These motions are each defined by a waypointed end-effector pose trajectory that is solved using the robot’s motion planner, with the planning solution then sent to the robot’s controllers. The pose trajectories are different between strategies, enabling the robot to manipulate the device in different ways. While the strategies tested in this work all use the same subtree design, this is not a limitation of the adaptive BT, and in general each strategy can use a completely different design, enabling greater variety between strategies.
Task-Specific Conditions. If currently manipulating the device, the robot checks whether the device is considered sufficiently manipulated. This condition is specified differently depending on the particular task; for example, when tightening a valve, the current joint efforts are used to estimate tightness, while when pressing a button, the joint efforts are used to estimate whether it has been pressed.
Strategy-Specific Conditions. If approaching, grasping or manipulating the device, the robot checks that its end-effector pose is within tolerance of the current pose trajectory. When manipulating the device, the robot checks that its joint efforts do not exceed the safe limits for the strategy.
The robot also checks that the device handle pose is still within safe limits for continuous manipulation using the strategy. This is important since a task like tightening a valve may require multiple revolutions of the handle, yet the robot may not be physically able to do this in one motion due it its own joint limits. Thus, this condition tells the robot when it should stop and re-grasp the device handle, preventing the robot from overextending itself during manipulation.
Recovery/Exit. Both subtrees use the same design; the robot retracts its end-effector from the device (if grasped) and then resets to its starting pose and joint configuration. The retraction is also defined by a waypointed end-effector pose trajectory and solved via motion planning.
@@ VI Implementation
Test Platform. We used an ANYbotics ANYmal C quadruped[38] with mounted Unitree Z1 robotic arm[39], equipped with a wrist-mounted Intel Realsense Depth Camera D345[40] for perception. ROS, the Robot Operating System, was used for the implementation of the test setup as well as communication with hardware [41].
@@
Behavior Tree. BT.CPP 4.0 was used to construct the adaptive BT. The ActionNode types specific to this work take the form of ROS action and service clients, implemented as derived classes of the StatefulActionNode from the BT.CPP core library; this node type supports asynchronous actions and is recommended for this use case.
Strategy Selection. The SelectStrategy node estimates the optimal manipulation strategy. For this work, the optimal strategy $a^{*}$ completes the task in the shortest time while being feasible given the effort data recorded during manipulation. This is equivalent to the following problem:
$a^{*}=\min_{a\in\mathcal{A}_{\mathrm{tsk}}}t_{\mathrm{est}}(a)\quad\text{s.t.}\quad T_{\mathrm{max},i}(a)<T_{\mathrm{lim},i}(a)\;\forall i$ (1)
where strategies $a$ are selected from the set of available task strategies $\mathcal{A}_{\mathrm{tsk}}$, $t_{\mathrm{est}}(a)$ is the estimated time for task completion, $T_{\mathrm{max},i}(a)$ are the largest recorded absolute effort values for each arm joint $i$, recorded from previous manipulation attempts using $a$, and $T_{\mathrm{lim},i}(a)$ are the maximum permitted absolute effort values for each joint $i$ under strategy $a$.
The effort data used for strategy selection is recorded from the arm joint efforts during each manipulation attempt, and stored in a database accessible to SelectStrategy.
@@
Motion Planning. This work uses a whole-body motion planner made using EXOTica, the EXtensible Optimisation Toolset [42]. The planner solves an EXOTica DynamicTimeIndexedShootingProblem, tracking pre-defined waypointed end-effector pose trajectories for the approach, grasp, manipulate and retract stages of the manipulation task for each strategy. The pose trajectories are defined relative to the pose of the device handle; in this way, the planner is generalizable to arbitrary device poses and handle angles (so long as robot joint limits permit the motion). In the case of planning failure, the planner action nodes return Failure, prompting the adaptive BT to retry the task.
@@
Control. This work uses the standard controllers provided with the ANYmal C and Unitree Z1. These controllers are by no means perfect, and in practice there was often non-negligible deviation between the true end-effector pose and that in the motion plan, especially when the arm was outstretched; to ensure safe operation, this deviation is checked by the condition monitor during manipulation. This is a major cause of the re-attempts that would occasionally occur during testing as seen in section VIII. Using a bespoke whole-body controller that compensates for both the motion of the base and arm would improve performance, but this is beyond the scope of this work; instead, we demonstrate how our adaptive BT can be used to overcome these problems by autonomously preempting failure and retrying.
@@
Perception. An Apriltag [43] was used to identify the pose of the device handle in an ‘unrotated’ state, and invariant template matching [44] was used to identify the handle angle (and thus the pose of the rotated handle).
VII Experiments
Needle Valve. The robot twists a Tameson NLS-012 needle valve[45] from unknown starting tightness until tight. Two strategies were provided; the faster, simpler ‘low-torque’ strategy grasps the handle head-on while the slower, more complex ‘high-torque’ strategy grasps it from the side. Fig. 4 depicts these strategies. The experiment is conducted using each strategy on its own, as well as using both strategies such that the BT can adaptively select between them.

### v3

@@ Abstract
Many manipulation tasks pose a challenge since they depend on non-visual environmental information that can only be determined after sustained physical interaction has already begun. This is particularly relevant for effort-sensitive, dynamics-dependent tasks such as tightening a valve. To perform these tasks safely and reliably, robots must be able to quickly adapt in response to unexpected changes during task execution, and should also learn from past experience to better inform future decisions. Humans can intuitively respond and adapt their manipulation strategy to suit such problems, but representing and implementing such behaviors for robots remains a challenge. In this work we show how this can be achieved within the framework of behavior trees. We present the adaptive behavior tree, a scalable and generalizable behavior tree design that enables a robot to quickly adapt to and learn from both visual and non-visual observations during task execution, preempting task failure or switching to a different manipulation strategy. The adaptive behavior tree selects the manipulation strategy that is predicted to optimize task performance, and learns from past experience to improve these predictions for future attempts. We test our approach on a variety of tasks commonly found in industry; the adaptive behavior tree demonstrates safety, robustness (100% success rate) and efficiency in task completion (up to 36% task speedup from the baseline).
@@ I Introduction
Service robots can revolutionize how entire service-oriented domains work, with potential applications including autonomous maintenance in – often remote – industrial facilities [1, 2], such as oil rigs [3], or care robots in domestic environments [4, 5, 6]. To complete tasks effectively, robots must use observations from the environment to construct and execute a suitable plan; a number of frameworks have been developed for task and motion planning (TAMP) for manipulation [7, 8, 9].
This work is motivated by effort-sensitive, dynamics-dependent manipulation tasks for articulated devices commonly found in industrial settings, for example twisting a valve until tightened to a specified torque, pushing an e-stop button until pressed, or firmly picking and placing a fragile object. Robots generally require highly controlled environments with known task parameters for reliable planning and execution of manipulation tasks; however, the tasks we are interested in pose a significant challenge since they depend on information that can only be determined after sustained physical interaction has already begun. For example, a robot cannot determine the initial tightness of a valve or stiffness of a button from visual data alone, and can only gather this information by beginning to interact with it. Even then, in the absence of an accurate model for the dynamic profile for the device, the robot cannot know for how long it must twist the valve, or how hard it must push the button, to complete the task. Even if provided a model, stiction from corrosion or unclear servicing may lead to uncertainty and deviation from standard assumptions. To solve these kinds of tasks, we argue that the robot must be able to instead decide on an initial plan with which to start the task and adapt to observations it makes during task execution. Since the tasks we consider may require the robot to apply a significant amount of effort to manipulate the device, adaptability is not only critical for task completion but also for safe robot operation; failure to respond to a sudden unexpected spike in effort can cause catastrophic failure and even damage to the robot or environment.
@@
We also consider the utility of a robot having access to different manipulation strategies for the same task. A strategy is the specific approach that a robot may take to carry out a task. For manipulation tasks, different strategies may involve grasping the object differently, or even using a tool to assist manipulation. Different strategies can have different trade-offs, for example speed, complexity, effort and power expenditure; Fig. 1 presents examples for tightening a valve. We seek a system that can smartly select and switch between known strategies, ideally optimizing for performance by learning from past manipulation attempts.
Figure 1: Different strategies for tightening a needle valve. (1a) Quickly grasping head-on means that all applied torque is generated purely from the wrist joint effort, producing a fast but ‘low-torque’ strategy. (1b) Grasping from the side requires more complex motion but enables the robot to use the arm as a wrench, producing a slow but ‘high-torque’ strategy.
We tackle both problems at the task planning level within the framework of behavior trees, exploiting their reactivity and modularity. The key contribution of our work is the adaptive behavior tree, which enables a robot to handle uncertainty in the environment and quickly adapt to unexpected safety-critical changes during task execution, preempting task failure or switching to a more optimal strategy based on data gathered over previous attempts. It comprises two stages; a selector that uses data recorded over previous attempts to predict the optimal strategy to deploy, and a monitor that checks a set of conditions concurrently to strategy execution, halting the current strategy if the task is complete or a safety condition is violated and re-attempting the task if necessary. The design is highly modular and generalizable, allowing for easy scaling to new strategies and tasks. We test the adaptive behavior tree on a diverse range of case studies representing complex, effort-intensive manipulation tasks from industry.
@@ II-A Adaptive Control for Robotics
It may be argued that our problem of manipulating articulated devices with unknown and changing dynamics intrinsics can be solved with an appropriate adaptive controller for each task. Adaptive control has existed in the literature for many years, and generally focuses on online parameter estimation and learning for uncertain dynamic systems. There exist many good surveys on the literature for adaptive control [10, 11]. The approach has been extended to robotic manipulation tasks, for example estimating model parameters for an uncertain end-effector payload [12, 13].
@@ II-B Behavior Trees for Robotics
There are many examples demonstrating how BTs can be used to coordinate manipulation tasks [18, 19, 20] and robot control [21, 22, 23]; Iovino et. al. provide a comprehensive survey of BTs for robotics [14]. BTs for robotics can be hand-coded [24, 25, 26], planned analytically [27, 28, 29, 30, 31, 32], learned from demonstration [33, 34, 35] or even generated using Large Language Models (LLMs) [36, 37].
@@ III Behavior Trees – An Overview
A Sequence propagates the tick to its children from left to right, returning Success if and only if all their children return Success. When it receives Running, it returns Running and ticks the same child on the next tick, and when it receives Failure, it returns Failure and resets the sequence progress. A ReactiveSequence is similar, but when it receives Running, it returns Running and also ticks all children up to the current child on the next tick; thus all prior children are still checked for Failure, enabling reactive behavior.
A Fallback (known as a Selector in other works) propagates the tick to its children from left to right until one of them returns Success, at which point the Fallback stops and returns Success. When it receives Running, it returns Running and ticks the same child on the next tick. It returns Failure only if all children return Failure. A ReactiveFallback is similar, but when it receives Running, it returns Running and also ticks all children up to the current child on the next tick; thus all prior children are still checked for Success, enabling reactive behavior (much like the ReactiveSequence).
@@
A DecoratorNode modifies the return status of a single child. A Precondition is a special type of DecoratorNode that ticks the child only if a precondition is met, or else returns either Success or Failure (marked by ‘S’ or ‘F’ respectively).
@@ IV Adaptive BT
Fig. 3 presents the adaptive BT, as well as a demonstration of how it can be integrated within a wider robot system. The adaptive BT is composed of two core stages; the strategy selector first decides which strategy to use for the next attempt, based on current perception data as well as data recorded over past attempts (such as joint efforts or time taken). The condition monitor then wraps around the subtree for the selected strategy, checking a set of task- and strategy-specific conditions concurrently to strategy execution. The adaptive BT also admits subtrees to coordinate robot setup prior to strategy selection, and for safely exiting or recovering from the current attempt in preparation for the next.
Figure 3: The adaptive BT, integrated within a wider robot system representing the test platform in this work. The colored lines represent communication lines for perception (blue), motion planning (green) and strategy selection (red).
@@
The condition checking is such that if the subtree of task-specific conditions returns Success, the task is considered to have been completed successfully, and if it returns Failure, the task is still ongoing. This follows the BT design principle of backward chaining [15], commanding the robot to keep executing the strategy until task completion conditions are met. Meanwhile, if the subtree of strategy-specific conditions returns Failure, a constraint has been violated during execution of the current strategy and the robot must halt the current attempt and try again, potentially selecting a different strategy. If it returns Success, no violations have occurred. This logic makes the condition monitor reactive and interpretable. If there are multiple task- or strategy-specific conditions to check, they should be placed in a ReactiveSequence to ensure proper operation of the adaptive BT. Each condition is ticked for every tick of the strategy subtree, allowing the robot to rapidly respond to new observations; thus the inherent reactivity of BTs is exploited. If a condition is not relevant to the current stage of the strategy subtree, it can be skipped by wrapping with an appropriate Precondition.
@@
The adaptive BT can easily be added as a subtree within a greater BT structure, and is thus readily applicable to existing BT-based robot task planning systems without major changes. The only required knowledge for adding the proposed design into a larger BT is that the adaptive BT must be repeatedly ticked while Running, and will return Success if the task was completed successfully and Failure otherwise.
The adaptive BT is designed to coordinate efficient selection, monitoring and switching of strategies, not the synthesis of the strategies themselves. The adaptive BT would synergize well with methods for the automatic synthesis of BTs [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37] to generate new strategies for it to use, but exploring this is beyond the scope of this work.
@@ V Adaptive BT for Manipulation
We now apply the adaptive BT to manipulation tasks; Fig. 4 presents the general subtrees for all strategies used in this work. We stress that these designs are just examples used to test the adaptive BT, with an emphasis on simplicity; the adaptive BT readily accepts new strategies for new tasks.
Figure 4: Subtrees for strategies used in this work. (4a) Manipulating an articulated device. (4b) Grasping an object. (4c) Moving an object to a goal.
Manipulating Articulated Devices. Fig. 4a presents general subtree structures for strategies that manipulate an articulated device, where task success occurs when the device is considered sufficiently manipulated (e.g. valve is tight). The strategies consist of approaching, grasping and manipulating the device (e.g. twist by one-half rotation), while checking safety conditions. Note that these same subtree structures are used for both the low-torque and high-torque strategies depicted in Fig. 1, and even the strategies for manipulating the different devices (e.g. tightening a valve vs pressing a button), though the action and condition nodes have different parameters (or even implementations) for each strategy.
Object Grasping. Fig. 4b presents general subtree structures for strategies that grasp an object, where task success occurs when the object is considered firmly grasped (e.g. by checking gripper effort). The strategies consist of approaching and grasping, while checking safety conditions.
Object Placing. Fig. 4c presents general subtree structures for strategies that move an (already grasped) object to a goal pose, where task success occurs if the strategy completes without preemption (thus no task-specific conditions). The strategies consist of moving and then releasing the object, while checking safety conditions.
Setup/Recovery/Exit. In this work, these are simply action nodes that position the robot to view the scene for strategy selection (for setup) or reset the robot pose (otherwise).
Strategy Selection. The SelectStrategy node estimates the optimal manipulation strategy to use; for this work, the optimal strategy $a^{*}$ for task state $s$ minimizes task duration $t_{\mathrm{est}}$ while not violating strategy-specific constraints (often corresponding to the safety conditions, e.g. maximum safe joint efforts). This is equivalent to the following problem:
$a^{*}(s)=\min_{a\in\mathcal{A}_{\mathrm{task}}}t_{\mathrm{est}}(a,s)\quad\text{subject to}\quad a\in\mathcal{A}_{\mathrm{safe}}(s),$
where strategies $a$ are selected from the set of available task-relevant strategies $\mathcal{A}_{\mathrm{task}}$, and $\mathcal{A}_{\mathrm{safe}}\subseteq\mathcal{A}_{\mathrm{task}}$ is the set of feasible strategies given the constraints. Both $t_{\mathrm{est}}$ and $\mathcal{A}_{\mathrm{safe}}$ are estimated from data recorded from previous attempts; if no data exists at $s$, we let $\mathcal{A}_{\mathrm{safe}}=\mathcal{A}_{\mathrm{task}}$, and calculate $t_{\mathrm{est}}$ from nominal durations of the strategy’s motion trajectories.
@@ VI Implementation
Test Platform. We used an ANYbotics ANYmal quadruped with mounted Unitree Z1 robotic arm, equipped with a wrist-mounted Intel Realsense Depth Camera D345 for perception. ROS, the Robot Operating System, was used for the implementation of the test setup as well as communication with hardware [39].
@@
Behavior Tree. BT.CPP 4.0 [38] was used to construct the adaptive BT. The ActionNode types specific to this work take the form of ROS action and service clients, implemented as derived classes of the StatefulActionNode from the BT.CPP core library; this node type supports asynchronous actions and is recommended for this use case.
@@
Motion Planning. This work uses a whole-body motion planner made using EXOTica, the EXtensible Optimisation Toolset [40]. The planner solves an EXOTica DynamicTimeIndexedShootingProblem, tracking pre-defined waypointed end-effector pose trajectories for the approach, grasp, manipulate and retract stages of the manipulation task for each strategy. The pose trajectories are defined relative to the pose of the object or device handle; in this way, the planner generalizes to arbitrary poses and handle angles (so long as robot joint limits permit the motion). In the case of planning failure, the planner action nodes return Failure, prompting the adaptive BT to retry the task.
@@
Control. This work uses the standard controllers provided with the ANYmal and Unitree Z1. The combined controller is by no means perfect, since the ANYmal controller has no awareness of the arm and cannot compensate for the induced wrench on the quadruped when the arm is outstretched. In practice, there was often non-negligible deviation between the true end-effector pose and that in the motion plan, especially when the arm was outstretched; to ensure safe operation, this deviation is checked by the condition monitor during manipulation. Using a bespoke whole-body controller that compensates for both the motion of the quadruped and arm would improve performance, but this is beyond the scope of this work; instead, we demonstrate how our adaptive BT can be used to overcome such problems at the task planning level by autonomously preempting failure and retrying.
@@
Perception. Apriltags [41] were used to identify the pose of objects as well as device handles in an ‘unrotated’ state, and invariant template matching [42] was used to identify handle angles and the pose of the rotated handles.
VII Case Studies
All device manipulation tasks use the general subtree structures depicted by Fig. 4a, albeit with task- and strategy-specific implementations of the action and condition nodes. Appendix A describes estimation of $s$ and $\mathcal{A}_{\mathrm{safe}}$ for all tasks.